OpenStudy (anonymous):

Factor this expression completely. 20w^2 - 45w + 10

OpenStudy (anonymous):

69

OpenStudy (anonymous):

not simplify, factor

OpenStudy (anonymous):

Do you know how to factor polynomials?

OpenStudy (the_fizicx99):

Well they don't have a variable gcf, however the do have a coefficient gcf. 10 = 5 * 2 20 = 5 * 2 * 2 45 = 5 * 9 The have a gcf of 5, So now divide - \[\frac{ 20w^2 - 45w + 10 }{ 5 } = ?\] Can you continue?

OpenStudy (anonymous):

YES! THANK YOU SO MUCH!!!